There are 21 animals on Farmer Cole's farm- all sheep and chickens. If the animals have a total of 56 legs, how many of each type of animal lives on his farm
Answer by checkley77(12844)

4s+2c=56
s+c=21 or s=21-c
4(21-c)+2c=56
84-4c+2c=56
-2c=56-84
-2c=-28
c=-28/-2
c=14 chickens.
s=21-14=7 sheep.
Proof:
4*7+2*14=56
28+28=56
56=56